Last night I found that the response of ITMX against the angle offsets were strage.
Eventually I found a loose connection at the feedthrough connectors of ITMX chamber.
So I pushed the connector hard, and then ITMX successfully became normal.
It looked like someone had accidentally kicked the cable during some works.
This bad connection had made unacceptable offsets in the OSEM readout, but now they seem fine.